# Exploring Neural Granger Causality with xLSTMs: Unveiling Temporal Dependencies in Complex Data

### v4

@@ Abstract
Causality in time series can be difficult to determine, especially in the presence of non-linear dependencies. The concept of Granger causality helps analyze potential relationships between variables, thereby offering a method to determine whether one time series can predict—Granger cause—future values of another. Although successful, Granger causal methods still struggle with capturing long-range relations between variables. To this end, we leverage the recently successful Extended Long Short-Term Memory (xLSTM) architecture and propose Granger causal xLSTMs (GC-xLSTM). It first enforces sparsity between the time series components by using a novel dynamic loss penalty on the initial projection. Specifically, we adaptively improve the model and identify sparsity candidates. Our joint optimization procedure then ensures that the Granger causal relations are recovered robustly. Our experimental evaluation on six diverse datasets demonstrates the overall efficacy of our proposed GC-xLSTM model.
@@ 1 Introduction
Finding cause and effect among and within a group of multivariate time series can lead to a better understanding of the dynamics of the involved individual time series. This can have an impact on critical research areas, such as finance [Masini et al., 2023], climate science [Mudelsee, 2019], and industrial applications [Strem et al., 2025]. Although recent efforts have been made to improve the interpretability of time series models [Ismail et al., 2020, Turbé et al., 2023], most methods are usually restricted to finding post-hoc interpretations and only focus on short-term dependencies.
@@
Many families of deep learning architectures have been explored for time series analysis over the years, such as multilayer perceptrons [Zeng et al., 2023, Das et al., 2023], recurrent neural networks (RNNs) [Hochreiter and Schmidhuber, 1997, Cho et al., 2014], convolutional neural networks [Wu et al., 2022, Wang et al., 2022], Transformers [Vaswani et al., 2017, Nie et al., 2023], state-space models [Wang et al., 2025], or mixing architectures [Wang et al., 2024]. Throughout this, recurrent models remained a natural choice for time series data since their direction of computation aligns well with the forward flow of time. This aligns particularly well with the goals of neural Granger causality. Furthermore, their inference runtime is typically linear in the sequence length at constant memory cost, making them much more efficient than, for instance, Transformers with quadratic runtimes and memory requirements, while remaining highly expressive. Recently, Beck et al. [2024] revisited recurrent models by borrowing insights gained from Transformers in many domains, specifically natural language processing. Their proposed Extended Long Short-Term Memory (xLSTM) model sparked a resurgence of interest in recurrent architectures for sequence modeling and have already proven highly suitable for time series forecasting [Kraus et al., 2024, Alharthi and Mahmood, 2024].
@@
Although most Granger causal machine learning methods assume linearity in time series as a fundamental prerequisite [Siggiridou and Kugiumtzis, 2015, Zhang et al., 2020], recent efforts capture non-linear dynamics in time series by using neural networks as the modeling choice instead of VARs [Tank et al., 2022, Löwe et al., 2022, Cheng et al., 2024]. Although successful, these non-linear methods require careful feature engineering to include time-based patterns and thus may not capture interactions between time series and external factors as effectively as xLSTMs, which can learn non-linear patterns and adapt to the non-stationary nature of time series data.
@@
We introduce GC-xLSTM, a novel method that leverages xLSTMs to uncover the GC relations in the presence of complex data, which inherently can have long-range dependencies. GC-xLSTM first enforces sparsity between the time series components by using a novel lasso penalty on the initial projection layer of the xLSTM. We learn a weight per time series and then adapt them to find the relevant variates for that step. Then, each time series component is modeled using a separate xLSTM model, enabling us to discover more interpretable GC relationships between the time series variables. After the forecast results by the individual xLSTM models, the important features are made more prominent, whereas the less important ones are diminished by a joint optimization technique, which includes using a novel reduction coefficient. Thus, the overall GC-xLSTM model can be trained end-to-end to uncover long-range Granger causal relations.
Key Contributions.   Our main research contributions can be summarized as follows: (1) We propose GC-xLSTM [^1] We make our code freely available at https://anonymous.4open.science/r/GC-xLSTM-A9FE., a novel model that can uncover Granger causal relations in non-linear time series. (2) Our novel algorithm jointly improves the forecasting model while adaptively enforcing strict sparsity. (3) We show that our proposed training regime increases the overall robustness of the underlying model, making it less sensitive to noise. (4) Our empirical evaluations demonstrate that GC-xLSTM can robustly discover Granger causal relations in the presence of complex simulated and real-world data.
@@
Outline.   We start by recalling preliminaries and reviewing related research to contextualize this work in the wider body of research on neural Granger causality in Section 2. This allows us to introduce GC-xLSTM in Section 3 and empirically evaluate in relation to other methods in Section 4. Finally, we conclude with an outlook to future work in Section 5.
@@ 2 Preliminaries and Related Work
We are interested in datasets of time series $\bm{S}\in\mathbb{R}^{V\times T}$ of $V$ variates with length $T$. Let $\bm{S}_{t}\in\mathbb{R}^{V}$ denote the value of $\bm{S}$ at time $t$. A variate $v$ (sometimes called a channel) can be any scalar measurement, such as the chlorophyll content of a plant or the spatial location of some object being tracked. Its value at time $t$ is $S_{v,t}\in\mathbb{R}$. The measurements are assumed to be carried out jointly at $T$ regularly spaced time steps. In forecasting, a model is presented with a time series of $C$ context steps $\bm{S}_{<t}$ before $t$, from which it shall predict the next value $\bm{S}_{t}\in\mathbb{R}^{V}$.
@@ 2.2 Neural Granger Causality
It would be very costly to train a total of $V^{2}$ models to test if each variate Granger causes any other variate and thus construct the entire Granger causal graph. To avoid this, we can train merely a single component-wise model $\mathcal{M}_{\bm{\theta},v}$ for each variate $v$ and inspect what inputs $w$ it is sensitive to. This can, for instance, be achieved by optimizing the predictive loss $\mathcal{L}_{\text{pred}}\left(\bm{\theta}\right)$ based on all model parameters $\bm{\theta}$ together with a regularizer $\Omega(\widehat{\bm{\theta}}_{v})$ that enforces sparse usage of the input features:
@@
where $\widehat{\bm{\theta}}_{v}$ are tunable parameters of the regularizer for variate $v$ and $\lambda\in\mathbb{R}_{+}$ is a hyperparameter to adjust the degree of sparsity. One such approach are cMLPs [Tank et al., 2022]; multilayer perceptrons where the first weight matrix $\bm{W}\in\mathbb{R}^{D\times V}$ projecting from $V$ features at each time step to $D$ hidden dimensions is regularized to encode a sparse selection of input features. $\Omega$ is instantiated as an $L^{2}$ norm of its columns: $\sum_{v=1}^{V}\left\lVert\bm{W}_{v}\right\rVert_{2}$. Note that $\bm{\theta}$ and $\widehat{\bm{\theta}}$ can overlap. Sparsity can then be extracted by binarizing the entries of $\bm{W}$ using a user-defined threshold $\tau$. This is necessary as the $L^{2}$ penalty tends only to shrink parameters to small values near zero, yet not clamp them sharply to it. This, however, allows subsequent layers to amplify the dampened signal again and still use it for forecasting. We avoid this disadvantage in GC-xLSTM by explicitly optimizing the feature extractor for strict sparsity. This more principled approach works without determining a sparsity threshold $\tau$. Previous work has explored both more regularizers [Tank et al., 2022] and different means to extract Granger causal relationships, including using feature attribution via explainability [Atashgahi et al., 2024] and interpretability [Marcinkevičs and Vogt, 2021]. Furthermore, several works have gone towards learning relevant representations that respect the underlying Granger causality [Xu et al., 2016, Varando et al., 2021, Dmochowski, 2023]. Zoroddu et al. [2024] present another approach where prior knowledge is encoded in the form of a noisy undirected graph, which aids the learning of Granger causality graphs. A more recent approach [Lin et al., 2024] employs Kolmogorov-Arnold networks [Liu et al., 2024] to learn the Granger causal relations between time series.
@@ 2.3 Extended Long Short-Term Memory (xLSTM)
Beck et al. [2024] propose two building blocks to build up xLSTM architectures: the sLSTM and mLSTM modules for vector-valued (multivariate) sequences. sLSTM cells improve upon classic LSTMs by exponential gating. For parallelizable training, mLSTM cells move from memory mixing between hidden states to an associative matrix memory instead. We will continue by recalling how sLSTM cells function since we found their mixing of consecutive memory states more effective in time series forecasting.
@@ 3.1 Overall Architecture
As Figure 1 shows, we estimate a pipeline of sparse feature selectors and xLSTM models to predict the multivariate time series. Eventually, this allows us to derive Granger causal dependencies from the selected features. Specifically, we learn for each variate $v$ a separate sparse projection and compute $\bm{x}_{v}=\bm{W}_{v}\bm{S}+\bm{b}_{v}$ shared across time. The matrix $\bm{W}_{v}\in\mathbb{R}^{D\times V}$ is shared across all lags of the time series for simplicity. We will lift this restriction in Section 4.2. Note that $\bm{b}$ does not affect the sparse use of inputs. We write $\bm{\phi}$ for the set of parameters $\bm{W}_{v}$ and $\bm{b}_{v}$ for all $v\in\mathcal{V}$. In addition to selecting dependencies, the sparse projection embeds the data into $D$-dimensional hidden space.
For Neural Granger Causality to successfully and faithfully extract the proper underlying dependencies, it is essential to employ models that can capture the complete set of dependencies. We, therefore, employ powerful deep-learning models with significantly higher capacity than the cMLPs and cLSTMs in prior work [Tank et al., 2022]. In particular, we instantiate the individual time series forecasters $\mathcal{M}_{\bm{\theta}_{v},v}$ with sLSTM blocks as introduced in Section 2.3. They can capture long-range dependencies in time series data, substantially enhancing the capabilities of traditional LSTMs in handling extended contexts. GC-xLSTM consists of $V$ sLSTM models, each modeling a different time series component. They are trained using established forecasting losses, such as the MSE loss $\mathcal{L}_{\text{pred}}\left(\bm{\phi},\bm{\theta}\right)$ from Eq. 1.
@@ 3.2 Optimizing for Strict Input Sparsity
(a) The optimization procedure for GC-xLSTM alternating between GD and proximal GD.
(b) The intuition of the compression step in 17. The matrix preserves columns with high norm important for Granger causality detection (dark) while suppressing others (light).
Figure 2: Optimization procedure and compression intuition for GC-xLSTM. Left: The alternating optimization procedure between GD and proximal GD. Right: Visual illustration of the compression mechanism, where important features are preserved while less relevant ones are reduced.
The purpose of the feature selector $\bm{W}$ is to only “pay attention” to as many variates as necessary for successful forecasting. A common approach to achieve this sparsity on the Granger causal relationships is via the lasso regularization explained in Section 2.2. We use a variation of the group lasso penalty [Yuan and Lin, 2006, Simon and Tibshirani, 2012] on the initial projection layer of GC-xLSTM as a structured sparsity-inducing penalty that encourages the selection of entire groups of variables, encoded as the columns of $\bm{W}_{v}$. Note that our penalty is different from adaptive lasso [Yuan and Lin, 2006], where the weights are not learned, but rather are treated as fixed heuristics. Note that standard gradient descent methods cannot directly optimize such a penalty due to its non-differentiability. We thus adaptively compress by learning a reduction coefficient $\bm{\alpha}_{v}\in\mathbb{R}^{V}$ that selects which of the $V$ columns of $\bm{W}_{v}$ are redundant. We perform this compression of $\bm{W}_{v}$ in a joint procedure with the general optimization of the forecasting model, as provided in Algorithm 1. Specifically, we perform two updates per optimization step for each of the variates, as Figure 2(a) depicts. Firstly, we optimize the projection weights $\bm{\phi}$, the reduction coefficients $\bm{\alpha}$, and the xLSTM parameters $\bm{\theta}$ using mini-batch gradient descent. This corresponds to 11 to 16 in Algorithm 1. This optimizes the following loss expected over the time series data $\bm{S}$:
@@
Note that we, crucially, only descend on the reduction coefficient $\bm{\alpha}$ in $\mathcal{L}_{\text{red}}$ and not on $\bm{W}_{v}$, as the stop-gradient $\operatorname{sg}(\cdot{})$ denotes. This sparsity optimization is instead performed by the second step in the procedure, shown in 17, where a proximal gradient descent step is taken to dynamically shrink $\bm{W}_{v}$ proportional to $\bm{\alpha}_{v}$. The compression update takes a descent step towards the gradient of
@@
followed by a soft thresholding. Intuitively, the $\mathcal{L}_{\text{red}}$ component of Eq. 3 keeps $\bm{W}_{v}$ fixed while learning $\bm{\alpha}_{v}$, and Eq. 4 keeps $\bm{\alpha}_{v}$ fixed in the proximal step to compress $\bm{W}_{v}$. Figure 2(b) depicts the intuition of the proximal gradient step and soft-thresholding in 17.
Details on learning the reduction loss $\mathcal{L}_{\text{red}}$.   It is worth briefly discussing the use of the logarithm in Eq. 3. It mainly gives more equal weight to the decreases in $\bm{W}_{v}$ column norms and encourages learning of better sparse Granger causal relations. It furthermore normalizes the gradient updates to $\bm{\alpha}_{v}$. Empirically, this loss engineering allowed training models that were significantly more robust to noise and changes to the sparsity hyperparameter $\lambda$. This was reflected by a more stable variable usage and predictive loss $\mathcal{L}_{\text{pred}}$.
Ensuring non-negativity of the reduction coefficients $\bm{\alpha}$.   For the proximal update step to be well-behaved, we need to ensure that $\bm{\alpha}_{v}$ results in a convex combination of column weights, i.e., that $\bm{\alpha}_{v}^{w}>0$ for all $w\in\mathcal{V}$ and $\bm{\alpha}_{v}^{T}\mathds{1}=1$. We achieve this by re-parameterizing it as $\bm{\alpha}_{v}=\mathrm{softmax}\left(\bm{\beta}_{v}\right)$, and learning $\bm{\beta}_{v}$ instead of $\bm{\alpha}_{v}$.
Intuitive dynamics of the gradient update step.   The weights $\bm{W}_{v}$ in the reduction loss $\mathcal{L}_{\text{red}}$ of Eq. 3 only serve to learn good reduction coefficients $\bm{\alpha}_{v}$, and are not optimized themselves. Deriving the gradient of the penalty term $\mathcal{L}_{\text{reg}}$ with respect to the underlying $\bm{\beta}$ provides a helpful intuition of the training dynamics:
$\frac{\partial}{\partial\beta_{v}^{w}}\sum_{w=1}^{V}\alpha_{v}^{w}\left\lVert\bm{W}_{v}^{w}\right\rVert_{2}=\frac{\partial}{\partial{\beta_{v}^{w}}}\sum_{w=1}^{V}\mathrm{softmax}(\bm{\beta}_{v})^{w}\left\lVert\bm{W}_{v}^{w}\right\rVert_{2}=\alpha_{v}^{w}\left(\left\lVert\bm{W}_{v}^{w}\right\rVert_{2}-\sum_{w=1}^{V}\alpha_{v}^{w}\left\lVert\bm{W}_{v}^{w}\right\rVert_{2}\right)$
@@
Practical considerations.   Furthermore, we perform staged optimization of $\bm{\alpha}$, which is initialized to a uniform distribution by setting all $\bm{\beta}=0$. We only start training $\bm{\alpha}$ after exploring the prediction loss and having obtained a reasonably compressed forecaster, which is controlled by the hyperparameter $K$ in Algorithm 1 (see 14). While we present the method with mini-batch gradient descent for conciseness, improved convergence can be achieved with more modern optimizers, such as Adam [Kingma and Ba, 2017].

### v5

@@ Abstract
Causality in time series can be challenging to determine, especially in the presence of non-linear dependencies. Granger causality helps analyze potential relationships between variables, thereby offering a method to determine whether one time series can predict—Granger cause—future values of another. Although successful, Granger causal methods still struggle with capturing long-range relations between variables. To this end, we leverage the recently successful Extended Long Short-Term Memory (xLSTM) architecture and propose Granger causal xLSTMs (GC-xLSTM). It first enforces sparsity between the time series components by using a novel dynamic loss penalty on the initial projection. Specifically, we adaptively improve the model and identify sparsity candidates. Our joint optimization procedure then ensures that the Granger causal relations are recovered robustly. Our experimental evaluation on six diverse datasets demonstrates the overall efficacy of GC-xLSTM.
@@ 1 Introduction
Finding cause and effect among and within a group of multivariate time series can lead to a better understanding of the dynamics of the involved time series. For instance, in computational neuroscience and medicine, discovering brain connectivity assists in better understanding natural cognition [Smith et al., 2011]. Discovering inter-dependencies between time series also has a critical impact on many other research areas, such as finance [Masini et al., 2023], climate science [Mudelsee, 2019], and industrial applications [Strem et al., 2025]. Although efforts have recently been made to improve the interpretability of time series models [Ismail et al., 2020, Turbé et al., 2023], most methods are restricted to finding post-hoc interpretations and only focus on short-term dependencies.
@@
Many families of deep learning architectures have been explored for time series analysis over the years, such as multilayer perceptrons [Zeng et al., 2023, Das et al., 2023], recurrent neural networks (RNNs) [Hochreiter and Schmidhuber, 1997, Cho et al., 2014], convolutional neural networks [Wu et al., 2022, Wang et al., 2022], Transformers [Vaswani et al., 2017, Nie et al., 2023], state-space models (SSMs) [Wang et al., 2025], or mixing architectures [Wang et al., 2024]. Throughout this, recurrent models remained a natural choice for time series data since their direction of computation aligns well with the forward flow of time. This aligns particularly well with the goals of neural Granger causality. Although SSMs are comparable in that regard, RNNs tend to offer more powerful forecasting capabilities since they deteriorate less when modeling long-term dependencies. Furthermore, their inference runtime is typically linear in the sequence length at constant memory cost, making them much more efficient than, for instance, Transformers with quadratic runtimes and memory requirements, while remaining highly expressive. Recently, Beck et al. [2024] revisited recurrent models by borrowing insights gained from Transformers in many domains, specifically natural language processing. Their proposed Extended Long Short-Term Memory (xLSTM) model sparked a resurgence of interest in recurrent architectures for sequence modeling and has already proven highly suitable for time series forecasting [Kraus et al., 2025, Alharthi and Mahmood, 2024].
@@
Although most Granger causal machine learning methods assume linearity in time series as a fundamental prerequisite [Siggiridou and Kugiumtzis, 2015, Zhang et al., 2020], recent efforts capture non-linear dynamics in time series by using neural networks as the modeling choice instead of VARs [Tank et al., 2022, Löwe et al., 2022, Cheng et al., 2024]. Although successful, these non-linear methods require careful feature engineering to include time-based patterns. Thus, they may not capture interactions between time series and external factors as effectively as xLSTMs, which can learn non-linear patterns and adapt to the non-stationary nature of time series data.
@@
We introduce GC-xLSTM, a novel method that leverages xLSTMs to uncover the GC relations in the presence of complex data, which inherently can have long-range dependencies. [^1] Code available at github.com/harpoonix/GC-xLSTM. GC-xLSTM first enforces sparsity between the time series components by using a novel lasso penalty on the initial projection layer of the xLSTM. We learn a weight per time series and then adapt them to find the relevant variates for that step. Then, each time series component is modeled using a separate xLSTM model, enabling us to discover interpretable GC relationships between the time series variables. After the forecast results by the individual xLSTM models, the important features are made more prominent, whereas the less important ones are diminished by a joint optimization technique, which includes using a novel reduction coefficient. Thus, the overall GC-xLSTM model can be trained end-to-end to uncover long-range Granger causal relations.
Our main research contributions can be summarized as follows:
  1. (i)
    We propose GC-xLSTM, a novel model that can uncover Granger causal relations in non-linear time series.
  2. (ii)
    Our novel algorithm jointly improves the forecasting model while adaptively enforcing strict sparsity.
  3. (iii)
    Our empirical evaluations demonstrate that GC-xLSTM can robustly discover Granger causal relations in the presence of complex simulated and real-world data.
Outline.
@@
We start by recalling preliminaries and reviewing related research to contextualize this work in the broader body of research on neural Granger causality in Section˜2. This allows us to introduce GC-xLSTM in Section˜3 and empirically evaluate in relation to other methods in Section˜4. Finally, we conclude with an outlook to future work in Section˜5.
@@ 2 Preliminaries and Related Work
We are interested in datasets of strictly stationary time series $\bm{S}\in\mathbb{R}^{V\times T}$ of $V$ variates with length $T$. Let $\bm{S}_{t}\in\mathbb{R}^{V}$ denote the value of $\bm{S}$ at time $t$. A variate $v$ (sometimes called a channel) can be any scalar measurement, such as the chlorophyll content of a plant or the spatial location of some object being tracked. Its value at time $t$ is $S_{v,t}\in\mathbb{R}$. The measurements are assumed to be carried out jointly at $T$ regularly spaced time steps. In forecasting, a model is presented with a time series of $C$ context steps $\bm{S}_{<t}$ before $t$, from which it shall predict the next value $\bm{S}_{t}\in\mathbb{R}^{V}$.
@@ 2.2 Neural Granger Causality
It would be very costly to train a total of $V^{2}$ models to test if each variate Granger causes any other variate and thus construct the entire Granger causal graph. To avoid this, we can train merely a single component-wise model $\mathcal{M}_{\bm{\theta},v}$ for each variate $v$ and inspect what inputs $w$ it is sensitive to. While this does not ablate models as in classical GC, it reduces the number of models to be trained from quadratic to linear in $V$. This can, for instance, be achieved by optimizing the predictive loss $\mathcal{L}_{\text{pred}}\left(\bm{\theta}\right)$ based on all model parameters $\bm{\theta}$ with a regularizer $\Omega(\widehat{\bm{\theta}}_{v})$ enforcing sparsity in the input features:
@@
where $\widehat{\bm{\theta}}_{v}$ are tunable parameters of the regularizer for variate $v$ and $\lambda\in\mathbb{R}_{+}$ is a hyperparameter to adjust the degree of sparsity. One such approach are cMLPs [Tank et al., 2022]; multilayer perceptrons where the first weight matrix $\bm{W}\in\mathbb{R}^{D\times V}$ projecting from $V$ features at each time step to $D$ hidden dimensions is regularized to encode a sparse selection of input features. $\Omega$ is instantiated as an $L^{2}$ norm of its columns: $\sum_{v=1}^{V}\left\lVert\bm{W}_{v}\right\rVert_{2}$. Note that $\bm{\theta}$ and $\widehat{\bm{\theta}}$ can overlap.
Sparsity can then be extracted by binarizing the entries of $\bm{W}$ using a user-defined threshold $\tau$. This is necessary as the $L^{2}$ penalty tends only to shrink parameters to small values near zero, yet not clamp them sharply to it. This, however, allows subsequent layers to amplify the dampened signal again and still use it for forecasting. We avoid this disadvantage in GC-xLSTM by explicitly optimizing the feature extractor for strict sparsity. This more principled approach works without determining a sparsity threshold $\tau$.
Previous work has explored both more regularizers [Tank et al., 2022] and different means to extract Granger causal relationships, including using feature attribution via explainability [Atashgahi et al., 2024] and interpretability [Marcinkevičs and Vogt, 2021]. Furthermore, several works have gone towards learning relevant representations that respect the underlying Granger causality [Xu et al., 2016, Varando et al., 2021, Dmochowski, 2023]. Zoroddu et al. [2024] present another approach where prior knowledge is encoded in the form of a noisy undirected graph, which aids the learning of Granger causality graphs. A more recent approach [Lin et al., 2024] employs Kolmogorov-Arnold networks [Liu et al., 2024] to learn the Granger causal relations between time series. For a discussion of when Granger causality implies (Pearlian) causality, we refer to Pettenuzzo and White [2011] and Das and Babadi [2023], allowing to connect to works such as the one of Rubenstein et al. [2017].
@@ 2.3 Extended Long Short-Term Memory (xLSTM)
Beck et al. [2024] propose two building blocks to build up xLSTM architectures: the sLSTM and mLSTM modules for vector-valued (i.e., multivariate) sequences. sLSTM cells improve upon classic LSTMs by exponential gating. For parallelizable training, mLSTM cells replace memory mixing between hidden states with an associative matrix memory. We will continue by recalling how sLSTM cells function since we found their memory mixing more effective in time series forecasting.
@@ 3.1 Overall Architecture
As Figure˜1 shows, we estimate a pipeline of sparse feature selectors and xLSTM models to predict the multivariate time series. Eventually, this allows us to derive Granger causal dependencies from the selected features. Specifically, we learn for each variate $v$ a separate sparse projection and compute $\bm{x}_{v}=\bm{W}_{v}\bm{S}+\bm{b}_{v}$ shared across time. The matrix $\bm{W}_{v}\in\mathbb{R}^{D\times V}$ is shared across all lags of the time series for simplicity. We will lift this restriction in Section˜4.2. Note that $\bm{b}$ does not affect the sparse use of inputs. We write $\bm{\phi}$ for the set of parameters $\bm{W}_{v}$ and $\bm{b}_{v}$ for all $v\in\mathcal{V}$. In addition to selecting dependencies, the sparse projection embeds the data into $D$-dimensional hidden space.
For Neural Granger Causality to successfully and faithfully extract the proper underlying dependencies, it is essential to employ models that can capture the complete set of dependencies. We, therefore, employ powerful deep-learning models with significantly higher capacity than the cMLPs and cLSTMs in prior work [Tank et al., 2022]. In particular, we instantiate the individual time series forecasters $\mathcal{M}_{\bm{\theta}_{v},v}$ with sLSTM blocks as introduced in Section˜2.3. They can capture long-range dependencies in time series data, substantially enhancing the capabilities of traditional LSTMs in handling extended contexts. GC-xLSTM consists of $V$ sLSTM models, each modeling a different time series component. They are trained using established forecasting losses, such as the MSE loss $\mathcal{L}_{\text{pred}}\left(\bm{\phi},\bm{\theta}\right)$ from Eq.˜1.
@@ 3.2 Optimizing for Strict Input Sparsity
(a) The optimization procedure for GC-xLSTM alternates between GD and proximal GD. A high value of $\alpha_{v}^{w}$ entails a strong compression of column $\bm{W}_{v}$, i.e. more sparsity in depending on variate $w$.
(b) Visual representation of the compression step in Algorithm˜1, ˜17. The matrix preserves columns with high norm that are important for Granger causality detection (dark) while suppressing others (light).
Figure 2: Optimization procedure and compression intuition for GC-xLSTM.
The purpose of the feature selector $\bm{W}$ is to only “pay attention” to as many variates as necessary for successful forecasting. A common approach to achieve this sparsity on the Granger causal relationships is via the lasso regularization explained in Section˜2.2. We use a variation of the group lasso penalty [Yuan and Lin, 2006, Simon and Tibshirani, 2012] on the initial projection layer of GC-xLSTM as a structured sparsity-inducing penalty that encourages the selection of entire groups of variables, encoded as the columns of $\bm{W}_{v}$. Note that our penalty differs from adaptive lasso [Yuan and Lin, 2006], where the weights are not learned but are treated as fixed heuristics. Note that standard gradient descent methods cannot optimize such a penalty directly due to its non-differentiability. We thus adaptively compress by learning a reduction coefficient $\bm{\alpha}_{v}\in\mathbb{R}^{V}$ that selects which of the $V$ columns of $\bm{W}_{v}$ are redundant. We perform this compression of $\bm{W}_{v}$ in a joint procedure with the general optimization of the forecasting model, as provided in Algorithm˜1. Specifically, we perform two updates per optimization step for each of the variates, as Figure˜2(a) depicts. Firstly, we optimize the projection weights $\bm{\phi}$, the reduction coefficients $\bm{\alpha}$, and the xLSTM parameters $\bm{\theta}$ using mini-batch gradient descent. This corresponds to 11 to 16 in Algorithm˜1. It optimizes the following loss expected over the time series data $\bm{S}$:
@@
Note that we, crucially, only descend on the reduction coefficient $\bm{\alpha}$ in $\mathcal{L}_{\text{red}}$ and not on $\bm{W}_{v}$, as the stop-gradient $\operatorname{sg}(\cdot{})$ denotes. This sparsity optimization is instead performed by the second step in the procedure, shown in ˜17, where a proximal gradient descent step dynamically shrinks $\bm{W}_{v}$ proportional to $\bm{\alpha}_{v}$. The compression update takes a descent step towards the gradient of
@@
followed by a soft thresholding. Intuitively, the $\mathcal{L}_{\text{red}}$ component of Eq.˜3 keeps $\bm{W}_{v}$ fixed while learning $\bm{\alpha}_{v}$, and Eq.˜4 keeps $\bm{\alpha}_{v}$ fixed in the proximal step to compress $\bm{W}_{v}$. Figure˜2(b) depicts the intuition of the proximal gradient step and soft-thresholding in ˜17.
Details on learning the reduction loss $\mathcal{L}_{\text{red}}$.
It is worth briefly discussing the use of the logarithm in Eq.˜3. It mainly gives more equal weight to the decreases in $\bm{W}_{v}$ column norms and encourages learning of better sparse Granger causal relations. It furthermore normalizes the gradient updates to $\bm{\alpha}_{v}$. Empirically, this loss engineering allowed training models that were significantly more robust to noise and changes to the sparsity hyperparameter $\lambda$. This was reflected by a more stable variable usage and predictive loss $\mathcal{L}_{\text{pred}}$.
Ensuring non-negativity of the reduction coefficients $\bm{\alpha}$.
For the proximal update step to be well-behaved, we need to ensure that $\bm{\alpha}_{v}$ results in a convex combination of column weights, i.e., that $\bm{\alpha}_{v}^{w}>0$ for all $w\in\mathcal{V}$ and $\bm{\alpha}_{v}^{T}\mathds{1}=1$. We achieve this by re-parameterizing it as $\bm{\alpha}_{v}=\mathrm{softmax}\left(\bm{\beta}_{v}\right)$, and learning $\bm{\beta}_{v}$ instead of $\bm{\alpha}_{v}$.
Intuitive dynamics of the gradient update step.
The weights $\bm{W}_{v}$ in the reduction loss $\mathcal{L}_{\text{red}}$ of Eq.˜3 only serve to learn good reduction coefficients $\bm{\alpha}_{v}$, and are not optimized themselves. Deriving the gradient of the penalty term $\mathcal{L}_{\text{reg}}$ with respect to the underlying $\bm{\beta}$ provides a helpful intuition of the training dynamics:
$\frac{\partial}{\partial\beta_{v}^{w}}\sum_{w=1}^{V}\alpha_{v}^{w}\left\lVert\bm{W}_{v}^{w}\right\rVert_{2}=\frac{\partial}{\partial{\beta_{v}^{w}}}\sum_{w=1}^{V}\mathrm{softmax}(\bm{\beta}_{v})^{w}\left\lVert\bm{W}_{v}^{w}\right\rVert_{2}=\alpha_{v}^{w}\hskip-3.0pt\left(\hskip-2.0pt\left\lVert\bm{W}_{v}^{w}\right\rVert_{2}-\sum_{w=1}^{V}\alpha_{v}^{w}\left\lVert\bm{W}_{v}^{w}\right\rVert_{2}\hskip-2.0pt\right)$
@@
Practical considerations.
Furthermore, we perform staged optimization of $\bm{\alpha}$, which is initialized to a uniform distribution by setting all $\bm{\beta}=0$. We only start training $\bm{\alpha}$ after exploring the prediction loss and having obtained a reasonably compressed forecaster, which is controlled by the hyperparameter $K$ in Algorithm˜1 (see ˜14). While we present the method with mini-batch gradient descent for conciseness, modern optimizers, such as Adam [Kingma and Ba, 2017], can further improve convergence.